how would you do this problem?
solve(find the x intercepts)
x to the second power-8x=-16

Answer by stanbon(75887)

solve(find the x intercepts)
x to the second power-8x=-16
------------------------------------
x^2 - 8x + 16 = 0
(x-4)(x-4) = 0
x = 4
(4,0) is the x-intercept.
=================================
Cheers,
Stan H.